OpenStudy (happykiddo):

The speed of a wave pulse on a string depends on the tension, F, in the string and the mass per unit length, μ, of the string. Tension has SI units of kg.m.s-2 and the mass per unit length has SI units of kg.m-1 What combination of F and μ must the speed of the wave be proportional to?

OpenStudy (happykiddo):

answer is( F / μ) i just don't know why

OpenStudy (anonymous):

what's u stand for in that regard?

OpenStudy (anonymous):

coefficient for what

OpenStudy (happykiddo):

its the symbol for Mu

OpenStudy (anonymous):

I don't know what that means

OpenStudy (anonymous):

so tell me

OpenStudy (anonymous):

Hey without you telling me the background info of the question I cannot hel you

OpenStudy (happykiddo):

sorry my computer battery turned off, Mu is mass/length string

OpenStudy (anonymous):

Basically the mas each unit of length consists of.

OpenStudy (anonymous):

Ok tenser the line greater the velocity at which the wave transfers from position A to position B is that what you are not quite on the spot

OpenStudy (anonymous):

Greater the mass of the line per unit of length harder it is for the pulse to traverse because it takes greater energy to move greater mass to transfer the pulse.

OpenStudy (happykiddo):

Thank you for the help : )

OpenStudy (happykiddo):

F/u was incorrect

OpenStudy (happykiddo):

its sqrt(T/u)

OpenStudy (anonymous):

but you still get the idea.

OpenStudy (anonymous):

this is a dimensional analysis problem. You actually have to write out the dimensions to figure out the equation. You're trying to combine the units of T and µ to give m/s, the unit of speed. Both T and µ have units of kg, but speed doesn't, so you know it they must be divided so the kg will cancel. \[\frac{ [T] }{ [\mu] }=\frac{ \frac{ kg-m }{ s^2 } }{ \frac{ kg }{ m } }=\frac{ m^2 }{ s^2 }\] The unit you want is m/s, so you have to take the square root. \[[v]=\sqrt{\frac{ m^2 }{ s^2 }}=\frac{ m }{ s }\] \[v \propto \sqrt{\frac{ T }{ \mu }}\]